Suppose the number of friends a randomly chosen Facebook member has is normally distributed with mean 300 and standard deviation 70.
A simple random sample of 50 members is taken. Determine the probability that the sample average number of friends is fewer than 310.

Answer by stanbon(75887)

Suppose the number of friends a randomly chosen Facebook member has is normally distributed with mean 300 and standard deviation 70.
A simple random sample of 50 members is taken. Determine the probability that the sample average number of friends is fewer than 310.
z(300) = (310-300)/(70/sqrt(50)) = 10/9.8999= 1.0102
-------------------------
P(x-bar < 310) = P(z < 1.0102) = 0.8438
------------------
Cheers,
Stan H.
------------------